The mean home price in Raleigh,
North Carolina, is $217,600. Assuming that the home
prices are normally distributed with a standard deviation
of $36,400, what is the probability that a randomly
selected home in Raleigh has a price below $200,000?
Below $150,000?

Answer by Boreal(15235)

z=(x-mean)/sd
=(200,000-217,600)/36400=-0.48
=(150,000-217600)/36400=-1.84
probability z<-0.48 is 0.3156
probability z<-1.84 is 0.0329